# On the observation of cosmic strings via gravitational-wave lensing

### v2

@@ V.2 Quantitative analysis of the lensing mismodelling bias
Figure 8 summarizes our findings. The left column shows the mismatch between lensed and unlensed waveforms as a function of impact parameter $y$ and time delay $t_{\Delta}$. The central column displays the SNR loss due to the consistency test, quantified by the ratio $\hat{\rho}/\rho$, where $\hat{\rho}$ is the reweighted SNR from Eq. (32), and $\rho$ is the matched-filter SNR obtained using an unlensed template bank. The right column shows the total SNR loss in the matched-filtering process, given by the ratio $\hat{\rho}/\rho_{\rm opt}$, where $\rho_{\rm opt}$ is the optimal SNR of the injected lensed signal. Each row in the figure corresponds to one of the three injections listed in Table 3, ordered by increasing SNR from top to bottom. To relate these results to the frequency of the beating pattern $f_{\Delta}$ (defined in Eq. (15)), combinations of $y$ and $t_{\Delta}$ that give frequencies between 20 Hz and 1000 Hz are indicated by dashed red lines (black in the rightmost column).
From Eq. (15) one can define a characteristic timescale at which the GO term becomes dominant. Using the time evolution of the GW frequency during the (adiabatic) inspiral phase of a compact binary coalescence [96],
@@
where $t_{\rm coal}$ represents the time of coalescence, we introduce a characteristic frequency that depends only on the chirp mass,
$f_{\mathrm{GO}}=\frac{1}{8\pi}\left(\frac{c^{3}}{G\mathcal{M}}\right)^{5/8}\,.$ (34)
This, in turn, defines a characteristic timescale,
$t_{\Delta}^{\mathrm{GO}}=\frac{4\pi}{y}\left(\frac{G}{c^{3}}\right)^{5/8}\mathcal{M}^{5/8}\approx\frac{0.012}{y}\,\times\left(\frac{\mathcal{M}}{M_{\odot}}\right)^{5/8}\,{\mathrm{s}}\,.$ (35)
@@
We also observe how the mismatch between the template bank and the detected signal impacts the SNR at different stages of the analysis. There are two distinct contributions to the overall SNR reduction: one arising from the waveform mismatch, which directly affects the matched-filter output, and another from the $\chi^{2}$ test. Interestingly enough, although both effects stem from the mismatch between the template and the signal, the $\chi^{2}$ test introduces a distinct pattern in the distribution of SNR downranking. Moreover, the penalization from the $\chi^{2}$ test increases with the SNR, in contrast to the mismatch-induced reduction, which is largely independent of the signal strength. The most significant SNR losses occur in the region of lensing parameters where the beating frequency $f_{\Delta}$ lies within the range [20,1000][20,1000]\,Hz, which overlaps well with the sensitivity band of the LVK detectors. In extreme cases, the total SNR reduction can reach up to $\sim 70\%$ of the injected SNR, with $\sim 60\%$ attributable to the $\chi^{2}$ consistency test alone. However, this reduction is less pronounced for injections with lower intrinsic SNR.
@@
Figure 9: Log odds ratio distributions as a function of lensing parameters: the time delay $t_{\Delta}$ and impact parameter $y$. The top row compares unlensed models with signals lensed by CSs. Positive values of the log odds ratio favor the CS lensing hypothesis, while negative values favor the unlensed model. The bottom row presents the same analysis, comparing CS lensing against lensing by a PML.
@@ V.3 Distinguishability of Lensed Signals
The mismatch between the signals in the template bank and the detected signal introduces a bias in matched filtering-based search. However, this mismatch also provides an opportunity to distinguish between different waveform models. To explore this, we employ Bayesian parameter estimation, which offers a framework for comparing the probabilities of competing hypotheses given a segment of strain data. The ratio of these probabilities is known as the odds ratio:
$\frac{p(\mathcal{H}_{1}|d)}{p(\mathcal{H}_{2}|d)}=\frac{p(d|\mathcal{H}_{1})}{p(d|\mathcal{H}_{2})}\times\frac{p(\mathcal{H}_{1})}{p(\mathcal{H}_{2})}\,.$ (36)
where the first term on the right-hand side is the Bayes factor (BF), which quantifies the ratio of evidences for the two hypotheses. It is typically computed using sampling methods that estimate the posterior distribution $p(\theta|d,\mathcal{H}_{i})$, allowing the evidence to be evaluated by marginalizing over the model parameters $\theta$:
$p(d|\mathcal{H})=\int p(d|\mathcal{H},\theta)p(\theta|\mathcal{H})\mathrm{d}\theta\,.$ (37)
In our analysis, we used the Bilby [100] library and the nessai [101] sampler for the BF calculation. The second term on the RHS of Eq. (36), known as the prior ratio, weights the BF by the relative prior probabilities of the two hypotheses.
In this work we consider the distinguishability between BBH waveforms lensed by CSs and (i) unlensed waveforms, and (ii) waveforms lensed by a PML. For the first case, assuming that the generation of a detectable BBH signal and its lensing are independent events, the prior ratio becomes:
@@
where $\mathcal{H}_{\rm U}$ denotes the unlensed hypothesis. The lensing probability can be estimated as:
@@
where $N_{\rm LS}$ is the number of detected lensed signals and $N_{\rm US}$ is the number of unlensed detected signals. The approximation assumes $N_{\rm LS}\ll N_{\rm US}$. Substituting $N_{\rm LS}$ with a central value of our calculated detection rate for CS-lensed signals (Appendix C) ($\sim 10^{-5}$) and $N_{\rm US}$ with the O4a detection rate (128 detections in 298 days, corresponding to $\sim 156.7$ events/year) [102], we obtain an estimate for the prior probability as $\log(p_{\rm CS})\approx-7.2^{-5.2}_{-9.2}$ depending on our assumption on the value of the $\Delta$ parameter.
For the comparison between CS-lensed and PML-lensed waveforms, the prior ratio is similarly given by:
@@
Using detection ratios, we estimate:
@@
To grossly estimate this prior ratio we will use the constraints on the strong lensing rate from [45], considering only the detection of single images, since these could already carry a microlensing imprint. By using these values we are assuming that strong lensing induced by galaxies or galaxy clusters also induce microlensing signatures. We note that this assumption produces an overestimation of the microlensing detection rate, as this might not be the case. For our analysis we will use, as a first order approximation, that the microlensing induced by galaxies and galaxy clusters can be modeled by the PML transmission factor and we will use a central value for the strong lensing rate of $\sim 10^{-3}$. We can estimate two bounds of the prior rate by looking at the two extreme cases: (i) the case with highest value of $\Delta$ and lower lensing rate, and (ii) the case with lowest $\Delta$ and highest lensing rate. This produces prior ratios in the following regime: $\log(\frac{p_{\rm CS}}{p_{\rm PML}})\approx-4.2^{-1.5}_{-6.5}$.
Bayesian model selection is highly sensitive to the SNR of the detected signal. Higher SNRs allow parameter estimation algorithms to extract more detailed features, improving model discrimination. Conversely, low-SNR signals may obscure key features that would give away information about the particular source of the signal. To quantify this effect, we compute the odds ratio for three BBH waveform injections, using the pre-lensing SNRs listed in Table 3. The used priors for the estimation of the BF are provided in Table 3.
Table 3: Priors used for the estimation of the BF. For the BBH parameters we used the default priors provided by Bilby, while the priors for the PML and CS lensing parameters were defined within the range where the beating pattern is visible in the frequency range of interest.
@@
$t_{\Delta}$ Uniform $5\cdot 10^{-4},1.3]$ ms
@@
Characteristic features of the CS transmission factor fall within the detector’s sensitivity band roughly when $t_{\Delta}>0.01$, as seen in the log-odds ratio plots, where amplification peaks and oscillations appear in the relevant frequency range. The behavior with respect to the impact parameter $y$ remains relatively consistent across the explored region.
@@ VI Conclusions
A central aspect of our work is the direct comparison between CS lensing and the standard PML model. The differences are fundamental and observationally significant. PML lensing is governed by a localized gravitational potential, admits the thin-lens approximation, and yields images determined by a lens equation. By contrast, CS lensing is topological, arising from the global conical structure of spacetime, which invalidates a strict thin-lens treatment. PMLs produce two images with magnifications and phase shifts set by the potential, whereas CSs generate two identical non-amplified images with no Morse phase shift. The parameter controlling wave effects in a PML is the distance-independent ratio $2R_{\rm S}/\lambda$, while for a CS it is $\mathcal{D}\Delta^{2}/\lambda$, which grows with the lens–observer distance $\mathcal{D}$. These differences lead to distinct waveform morphologies: microlensing appears as a beating pattern from overlapping images, while strong lensing yields two time-shifted replicas with delay $\Delta t_{21}=2y\,t_{\Delta}$. We derived a detectability bound on the string tension, Eq. (26), which sets a benchmark for searches.
@@
We also quantified the selection bias in searching for CS-lensed signals with unlensed template banks. The frequency modulation induced by $F(f)$ reduces the matched-filter SNR, with the largest losses occurring when the beating frequency $f_{\Delta}$ falls within the LVK band (202010001000 Hz). The $\chi^{2}$ signal-consistency test introduces an additional downranking, contributing up to $\sim 60\%$ of the SNR loss in extreme cases. The mismatch peaks at intermediate frequencies where detector sensitivity is greatest, amplifying the impact of lensing. Finally, Bayesian model selection shows that CS-lensed signals are distinguishable from both unlensed and PML-lensed hypotheses across a wide range of parameters. This distinctiveness, rooted in the characteristic phase and amplitude structure encoded in $F$, demonstrates that targeted searches can successfully identify CS lensing. It would also be interesting to explore the overlap between CS lensing and the phenomenological search for single distorted signals [77], which similarly produces beating patterns like those of PML and CS lensing. Assessing the distinguishability of CS-lensed signals in that broader context could provide further insight into their observational signatures. While this is beyond the scope of the present work, it represents a natural direction for future investigation. Overall, our framework offers a starting point for dedicated searches, enhancing the prospects of discovering cosmic strings through gravitational-wave lensing observations and opening a new window onto high-energy physics.
@@ Acknowledgements.
The authors are grateful to Mick Wright for his helpful comments on the manuscript. This work was supported by the Spanish Ministry of Science and Innovation through grants PID2021-125485NB-C21, PID2021-125485NB-C22, PID2024-159689NB-C21, and PID2024-159689NB-C22 funded by MCIN/AEI/10.13039/501100011033 and the European Regional Development Fund (ERDF), ‘A way of making Europe’. O.B. acknowledges support from the Agència de Gestió d’Ajuts Universitaris i de Recerca (AGAUR), Generalitat de Catalunya (grant SGR-2021-01069). N.V. and J.A.F. are supported by the Generalitat Valenciana (grant CIPROM/2022/49). J.A.F. acknowledges further support from the European Horizon Europe staff exchange (SE) programme HORIZON-MSCA2021-SE-01 Grant No. NewFunFiCO-101086251. The authors gratefully acknowledge the computer resources provided by the Nyx cluster at the ICCUB and by Artemisa, as well as the technical support from the Instituto de Física Corpuscular (IFIC, CSIC-UV). Artemisa is co-funded by the European Union through the 2014-2020 ERDF Operational Programme of the Comunitat Valenciana under project IDIFEDER/2018/048. This material is based upon work supported by NSF’s LIGO Laboratory which is a major facility fully funded by the National Science Foundation.

### v3

@@ V.2 Quantitative analysis of the lensing mismodeling bias
Figure 8 summarizes our findings. The left column shows the mismatch between lensed and unlensed waveforms as a function of impact parameter $y$ and time delay $t_{\Delta}$. The central column displays the SNR loss due to the consistency test, quantified by the ratio $\hat{\rho}/\rho$, where $\hat{\rho}$ is the reweighted SNR from Eq. (32), and $\rho$ is the matched-filter SNR obtained using an unlensed template bank. The right column shows the total SNR loss in the matched-filtering process, given by the ratio $\hat{\rho}/\rho_{\rm opt}$, where $\rho_{\rm opt}$ is the optimal SNR of the injected lensed signal. Each row in the figure corresponds to one of the three injections listed in Table 2, ordered by increasing SNR from top to bottom. To relate these results to the sensitivity band of the detectors, the relevant frequency range is indicated in Fig. 8 (see the figure caption for details).
By inspecting Eqs. (11) and (15), one can define a critical value for the characteristic time delay, $t_{\Delta}$, at which the GO term begins to dominate over the diffraction term. We denote this critical value as $t_{\Delta}^{\mathrm{GO}}$. Using the time evolution of the GW frequency during the (adiabatic) inspiral phase of a compact binary coalescence [97],
@@
where $t_{\rm coal}$ represents the time of coalescence, a reference frequency depends only on the chirp mass,
$f^{*}=\frac{1}{8\pi}\left(\frac{c^{3}}{G\mathcal{M}}\right)^{5/8}\,.$ (34)
When the characteristic frequency $f_{\Delta}$, given by Eq. (15), falls below $f^{*}$, the diffraction effects become negligible. This defines a critical value for $f_{\Delta}$,
$f_{\Delta}^{\mathrm{GO}}=\frac{1}{8\pi}\left(\frac{c^{3}}{G\mathcal{M}}\right)^{5/8}=\frac{1}{2t_{\Delta}^{\mathrm{GO}}y},$ (35)
which, in turn, defines the critical GO time delay mentioned above,
$t_{\Delta}^{\mathrm{GO}}=\frac{4\pi}{y}\left(\frac{G}{c^{3}}\right)^{5/8}\mathcal{M}^{5/8}\;\approx\;\frac{0.012\,{\mathrm{s}}}{y}\,\left(\frac{\mathcal{M}}{M_{\odot}}\right)^{5/8}.$ (36)
@@
We also observe how the mismatch between the template bank and the detected signal impacts the SNR at different stages of the analysis. There are two distinct contributions to the overall SNR reduction: one arising from the waveform mismatch, which directly affects the matched-filter output and another from the $\chi^{2}$ test. Interestingly enough, although both effects stem from the mismatch between the template and the signal, the $\chi^{2}$ test introduces a distinct pattern in the distribution of SNR downranking. Moreover, the penalization from the $\chi^{2}$ test increases with the SNR, in contrast to the mismatch-induced reduction, which is largely independent of the signal strength. The most significant SNR losses occur in the region of lensing parameters where the beating frequency $f_{\Delta}$ lies within the range [20,1000][20,1000]\,Hz, which overlaps well with the sensitivity band of the LVK detectors. In extreme cases, the total SNR reduction can reach up to $\sim 70\%$ of the injected SNR, with $\sim 60\%$ attributable to the $\chi^{2}$ consistency test alone. However, this reduction is less pronounced for injections with lower intrinsic SNR.
@@
[…]
@@ V.3 Distinguishability of Lensed Signals
The mismatch between the signals in the template bank and the detected signal introduces a bias in matched filtering-based searches. However, this mismatch also provides an opportunity to distinguish between different waveform models. To explore this, we employ Bayesian parameter estimation, which offers a framework for comparing the probabilities of competing hypotheses given a segment of strain data. The ratio of these probabilities is known as the odds ratio:
$\frac{p(\mathcal{H}_{1}|d)}{p(\mathcal{H}_{2}|d)}=\frac{p(d|\mathcal{H}_{1})}{p(d|\mathcal{H}_{2})}\times\frac{p(\mathcal{H}_{1})}{p(\mathcal{H}_{2})}\,$, (37)
where $\mathcal{H}_{i}$ represents the hypothesis under exploration, which can be either having a waveform lensed by a CS, by a PML or unlensed, $d$ refers to the observed data and $p$ represents the corresponding conditional probability. The first term on the right-hand side is the Bayes factor (BF), which quantifies the ratio of evidences for the two hypotheses. It is typically computed using sampling methods that estimate the posterior distribution $p(\theta|d,\mathcal{H}_{i})$, allowing the evidence to be evaluated by marginalizing over the model parameters $\theta$:
$p(d|\mathcal{H}_{i})=\int p(d|\mathcal{H}_{i},\theta)p(\theta|\mathcal{H}_{i})\mathrm{d}\theta\,.$ (38)
In our analysis, we used the Bilby [101] library and the nessai [102] sampler for the BF calculation. These libraries implement algorithms for estimating the posterior probability distribution $p(\theta|d)$ using nested sampling. The inference relies on a probabilistic description of the detector noise, typically assumed to be Gaussian. This approximation is accurate for sufficiently short, glitch-free signals, for which non-stationarity in the noise can be neglected. In our case, the noise is generated directly from the detector sensitivity curve and is therefore exactly Gaussian by construction.
Given this noise model, we can define a likelihood function $\mathcal{L}(d|\theta)$. Combined with specified prior distributions on the waveform parameters, the likelihood enables computation of the evidence (or marginal likelihood) by integrating the likelihood over the prior volume. Nested sampling performs this calculation by drawing parameter samples from the prior and iteratively replacing the lowest-likelihood sample with a new one of higher likelihood. This procedure transforms the multidimensional evidence integral into a one-dimensional integral over the decreasing sequence of constrained prior volumes. As higher-likelihood regions are explored, the evidence estimate converges, providing a principled termination criterion.
The nessai framework accelerates this process by learning the distribution of the region of parameter space with likelihood above the current threshold. It achieves this using normalizing flows—neural-network-based models capable of representing complex, non-Gaussian distributions—which greatly improve sampling efficiency and reduce the computational cost of nested sampling.
Once the sampler has converged, we obtain both the posterior distribution and a robust estimate of the evidence for the hypothesis. These evidence values become the numerator and denominator of the Bayes factor, thereby quantifying how well each waveform model explains the observed data. The second term on the RHS of Eq. (37), known as the prior ratio, weights the BF by the relative prior probabilities of the two hypotheses.
In this work we consider the distinguishability between BBH waveforms lensed by CSs and (i) unlensed waveforms, and (ii) waveforms lensed by a PML. For the first case, assuming that the generation of a detectable BBH signal and its lensing are independent events, the prior ratio becomes
@@
where $\mathcal{H}_{\rm U}$ denotes the unlensed hypothesis. The lensing probability can be estimated as
@@
where $N_{\rm LS}$ is the number of detected lensed signals per year and $N_{\rm US}$ is the number of unlensed detected signals per year. The approximation assumes $N_{\rm LS}\ll N_{\rm US}$. Substituting $N_{\rm LS}$ with a central value of our calculated detection rate for CS-lensed signals (Appendix C) ($\sim 10^{-5}$) and $N_{\rm US}$ with the O4a detection rate (128 detections in 298 days, corresponding to $\sim 156.7$ events/year) [103], we obtain an estimate for the prior probability as $\log(p_{\rm CS})\approx-7.2^{-5.2}_{-9.2}$ depending on our assumption on the value of the $\Delta$ parameter.
For the comparison between CS-lensed and PML-lensed waveforms, the prior ratio is similarly given by
@@
Using detection ratios, we estimate
@@
To grossly estimate this prior ratio we will use the constraints on the strong lensing rate from [46], considering only the detection of single images, since these could already carry a microlensing imprint. By using these values we are assuming that strong lensing induced by galaxies or galaxy clusters also induce microlensing signatures. We note that this assumption produces an overestimation of the microlensing detection rate, as this might not be the case. For our analysis we will use, as a first-order approximation, that the microlensing induced by galaxies and galaxy clusters can be modeled by the PML transmission factor and we will use a central value for the strong lensing rate of $\sim 10^{-3}$. We can estimate two bounds of the prior rate by looking at the two extreme cases: (i) the case with highest value of $\Delta$ and lower lensing rate and (ii) the case with lowest $\Delta$ and highest lensing rate. This produces prior ratios in the following regime: $\log(\frac{p_{\rm CS}}{p_{\rm PML}})\approx-4.2^{-1.5}_{-6.5}$.
Bayesian model selection is highly sensitive to the SNR of the detected signal. Higher SNRs allow parameter estimation algorithms to extract more detailed features, improving model discrimination. Conversely, low-SNR signals may obscure key features that would give away information about the particular source of the signal. To quantify this effect, we compute the odds ratio for three BBH waveform injections, using the prelensing SNRs listed in Table 2. The used priors for the estimation of the BF are provided in Table 3.
Table 3: Priors used for the estimation of the Bayes factor. For the BBH parameters we used the default priors provided by Bilby, while the priors for the PML and CS lensing parameters were defined within the range where the beating pattern is visible in the frequency range of interest.
@@
$t_{\Delta}$ Uniform [^7] The parameter $\Delta$ is not sampled independently, but is derived from $t_{\Delta}$ via Eq. (16), yielding $\Delta\in[7\times 10^{-11},\,3.5\times 10^{-9}]$ for $\chi_{L}=1\,$Gpc. We note that $\Delta\propto\chi_{L}^{-1/2}$ and thus decreases with increasing $\chi_{L}$. $5\cdot 10^{-4},1.3]\,$s
@@
Characteristic features of the CS transmission factor fall within the detector’s sensitivity band roughly when $t_{\Delta}>0.01\,$s, as seen in the log-odds ratio plots, where amplification peaks and oscillations appear in the relevant frequency range. The behavior with respect to the impact parameter $y$ remains relatively consistent across the explored region.
@@ VI Conclusions
A central aspect of our work is the direct comparison between CS lensing and the standard PML model. The differences are fundamental and observationally significant. PML lensing is governed by a localized gravitational potential, admits the thin-lens approximation, and yields images determined by a lens equation. By contrast, CS lensing is topological, arising from the global conical structure of spacetime, which invalidates a strict thin-lens treatment. PMLs produce two images with magnifications and phase shifts set by the potential, whereas CSs generate two identical non-amplified images with no Morse phase shift. The parameter controlling wave effects in a PML is the distance-independent ratio $2R_{\rm S}/\lambda$, while for a CS it is $\mathcal{D}\Delta^{2}/\lambda$, which grows with the lens-observer distance $\mathcal{D}$. These differences lead to distinct waveform morphologies: microlensing appears as a beating pattern from overlapping images, while strong lensing yields two time-shifted replicas with delay $\Delta t_{21}=2y\,t_{\Delta}$. We derived a detectability bound on the string tension, Eq. (26), which sets a benchmark for searches.
@@
We also quantified the selection bias in searching for CS-lensed signals with unlensed template banks. The frequency modulation induced by $F(f)$ reduces the matched-filter SNR, with the largest losses occurring when the beating frequency $f_{\Delta}$ falls within the LVK band (202010001000 Hz). The $\chi^{2}$ signal-consistency test introduces an additional downranking, contributing up to $\sim 60\%$ of the SNR loss in extreme cases. The mismatch peaks at intermediate frequencies where detector sensitivity is greatest, amplifying the impact of lensing. Finally, Bayesian model selection shows that CS-lensed signals are distinguishable from both unlensed and PML-lensed hypotheses across a wide range of parameters. This distinctiveness, rooted in the characteristic phase and amplitude structure encoded in $F$, demonstrates that targeted searches can successfully identify CS lensing. It would also be interesting to explore the overlap between CS lensing and the phenomenological search for single distorted signals [78], which similarly produces beating patterns like those of PML and CS lensing. Assessing the distinguishability of CS-lensed signals in that broader context could provide further insight into their observational signatures. While this is beyond the scope of the present work, it represents a natural direction for future investigation.
An important aspect of our analysis concerns the role of the string tension $\Delta$ and the lens distance $\chi_{L}$ in shaping observational signatures and their interplay with source parameter degeneracies. The lensing time delay $t_{\Delta}\propto\chi_{L}\Delta^{2}$ [Eq. (16)] directly controls the frequency scale of interference fringes, which can mimic intrinsic features of the binary waveform—such as spin-induced modulations or eccentricity—if not properly modeled. This highlights the need for dedicated template families that incorporate cosmic string lensing, as standard unlensed templates may misattribute lensing-induced modulations to astrophysical parameters.
Finally, Eqs. (25)–(27) quantify the detectability threshold for lensing signatures. Enforcing $f_{00}<f_{+}$ imposes a lower bound on the time delay $t_{\Delta}$, and hence on the string tension $\Delta$, which scales as
$\Delta_{\min}\propto(\chi_{L}f_{+})^{-1/2}.$ (43)
This dependence on the product of lens distance and detector frequency band introduces an important trade-off: while low-frequency detectors such as LISA require larger tensions for nearby lenses, this requirement can be partially compensated by cosmological distances, where $\chi_{L}$ is large and the time delay grows accordingly. For LVK ($f_{+}\sim 10^{3}$ Hz), the threshold is $\Delta\gtrsim 10^{-10}$ for lenses at $\sim 100$ Mpc, whereas next-generation ground-based detectors with higher $f_{+}$ will probe even smaller tensions. Conversely, LISA operating at millihertz frequencies would need $\Delta\gtrsim 10^{-8}$ for nearby lenses, but could detect weaker strings if they lie at gigaparsec scales. This frequency-distance interplay underscores the complementarity between detector bands: terrestrial interferometers are optimal for lensing by weak strings at moderate distances, while space-based missions target strong strings or distant lenses, alongside their sensitivity to direct GW emission from string networks.
Overall, our framework offers a starting point for dedicated searches, enhancing the prospects of discovering cosmic strings through gravitational-wave lensing observations and opening a new window onto high-energy physics.
@@ Acknowledgements.
The authors are grateful to Mick Wright for his helpful comments on the manuscript. This work was supported by the Spanish Ministry of Science and Innovation through grants PID2021-125485NB-C21, PID2021-125485NB-C22, PID2024-159689NB-C21, and PID2024-159689NB-C22 funded by MCIN/AEI/10.13039/501100011033 and the European Regional Development Fund (ERDF), ‘A way of making Europe’. O.B. acknowledges financial support from grant CEX2024-001451-M, funded by MICIU/AEI/10.13039/501100011033, and support from the Agència de Gestió d’Ajuts Universitaris i de Recerca (AGAUR), Generalitat de Catalunya (grant SGR-2021-01069). N.V. and J.A.F. are supported by the Generalitat Valenciana (grant CIPROM/2022/49). J.A.F. acknowledges further support from the European Horizon Europe staff exchange (SE) programme HORIZON-MSCA2021-SE-01 Grant No. NewFunFiCO-101086251. The authors gratefully acknowledge the computer resources provided by the Nyx cluster at the ICCUB and by Artemisa, as well as the technical support from the Instituto de Física Corpuscular (IFIC, CSIC-UV). Artemisa is cofunded by the European Union through the 2014-2020 ERDF Operational Programme of the Comunitat Valenciana under project IDIFEDER/2018/048. This material is based upon work supported by NSF’s LIGO Laboratory which is a major facility fully funded by the National Science Foundation.